# Space-based mm/mg-scale Laser Interferometer for Quantum Gravity

### v1

@@ Abstract
The experimental verification of the quantum nature of gravity represents a milestone in quantum gravity research. Recently, interest has grown for testing it via gravitationally induced entanglement (GIE). Here, we propose a space-based interferometer inspired by the LISA Pathfinder (LPF). Unlike the LPF, our design employs two smaller gold–platinum test masses, each weighing the milligram scale, surrounded by a shield below 2 K, and positioned side by side with a millimeter scale separation. This configuration enables the detection of GIE through simultaneous measurements of differential and common-mode motions. We simulate quantum measurements of these modes taking into account gas damping, black-body radiation, and cosmic-ray collisions to estimate the integration time for GIE detection. Our results show that GIE can be demonstrated with a few modifications to the LPF setup.
@@
Here, we propose a cavity-free laser interferometer to optimize the second approach in optomechanical systems, in contrast to previous cavity-based studies [19, 21, 14, 17, 23]. Our proposed experiment closely resembles the space-based interferometer LISA Pathfinder (LPF) [24], which marked the beginning of a new era in interferometric experiments; however, it differs in the following four key aspects: (1) The test masses are reduced to the milligram scale to suppress decoherence caused by blackbody radiation; (2) The test masses are placed close to enable the generation of gravitationally induced entanglement (GIE); (3) Both differential and common-mode motions are measured to enable the detection of GIE via Kalman filtering; and (4) The masses are surrounded by a shield below 2 K, and both modes of motion are further cooled via feedback using a high-pass filter. Using the space environment, the resonance frequency of the test mass can be reduced below 1 mHz free from ground vibrations, making it possible to observe the GIE within the total integration time of about 200 days.
Figure 1: Experimental setup. The differential and common-mode of test mass 1 (TM1) and test mass 2 (TM2) are measured by laser interferometry with a wavelength of $1064\ \mathrm{nm}$.
@@
where $Y_{\pm}$ represent the phase quadratures, $\bm{C}_{\pm}=(-\alpha_{\pm},0,0)$ denote the optomechanical coupling constants, and $\bm{x}_{\pm}=(q_{\pm},p_{\pm},r_{\pm})^{T}$ denote the position, momentum, and auxiliary variables used to describe feedback cooling (for details see Appendix A). The terms $y^{\rm in}_{\pm}$ represent vacuum (white) noise, with variances given by $\langle(y_{\pm}^{\rm in})^{2}\rangle=1$. In addition, we introduce vacuum noise $x^{\rm in}_{\pm}$ for the amplitude quadratures. Throughout this work, the subscripts $+$ and -- refer to the common mode and the differential mode, respectively. The optomechanical coupling constants $\alpha_{\pm}$ are given by $\sqrt{16\omega_{c}P_{\rm in}/(m\Omega_{\pm}c^{2}})$, where $\omega_{c}$ is the laser frequency and $P_{\rm in}$ the incident laser power, $m$ the mass of each mirror, $\Omega_{+}=\Omega$ and $\Omega_{-}=\Omega\sqrt{1-\delta}$ the resonance frequency of the common and differential modes. Here, $\delta=4Gm/(\Omega^{2}L^{3})$ denotes the gravitational coupling between the two mirrors, the resonance frequency for each mirror is denoted as $\Omega$, and $L$ represents the average distance between the mirrors. In this paper, we assume $m/L^{3}=\rho\Lambda$, where $\rho$ is the mass density of the mirror and $\Lambda$ is the factor determined by the shape and configuration of the mirrors [21].
@@
Here, $\bm{\mathcal{V}}$ is the conditional covariance matrix of the individual test masses (see Appendix B) and $\Sigma=\det\bm{\mathcal{V}}_{1}+\det\bm{\mathcal{V}}_{2}-2\det\bm{\mathcal{V}}_{12}$, where $\bm{\mathcal{V}}_{1}$, $\bm{\mathcal{V}}_{2}$ are the covariance matrices of TM1 and TM2 normalized by the frequency $\Omega$, respectively. $\bm{\mathcal{V}}_{12}$ represents the gravity-induced correlation matrix between individual masses. According to the separability condition for two-mode Gaussian states, the systems are entangled if and only if $E_{N}>0$ [27]. To verify whether quantum entanglement is present from Eq. (5), it is sufficient to confirm $\Sigma-\text{det}\bm{\mathcal{V}}>1$ for $\Sigma>2$.
The colored region in Fig. 2 shows the region where GIE is generated. Its generation rate must be faster than the thermal decoherence rate, as shown below:
@@
From the above inequality, the crucial condition for the GIE can be written as
${\left({\Gamma T/2\pi\over 10^{-18}{\ \rm Hz\cdot K}}\right)}{\left({\rho\over 20{\ \rm g/cm^{3}}}\right)}^{-1}\left({\Lambda\over 2}\right)^{-1}<2.4,$
where $T$ represents the temperature of the test masses. The required $\Gamma T/(2\pi)$ is about $10^{-18}\,\mathrm{Hz\cdot K}$.
Figure 2: Contour plot of $E_{N}$ in the $\alpha_{+}$-$\Gamma T/(2\pi)$ plane. The green dashed line indicates the threshold for generating the GIE in Eq. (6), the red curve shows the boundary where the relaxation time due to the Kalman filter is equal to half the thermal decoherence time (see Appendix A). The star symbol corresponds to the parameters used in the simulation.
@@
Finally, the acceleration noise due to cosmic-ray collisions is calculated using Geant4 [30] and a cosmic ray flux model [31]. By removing the effects of large events caused by cosmic-ray collisions through data processing and control, the required shielding thickness to sufficiently reduce the noise can be limited to about 1 m for 18 mg test masses. The duty cycle — defined as the fraction of time during which the data is usable for analysis — is about 56% (for details see Appendix C).
@@
In the simulation, we set $\Omega_{+}/(2\pi)=7.7\times 10^{-4}\ \mathrm{Hz}$, $\alpha_{+}=0.081\ \mathrm{Hz^{1/2}}$, $\Gamma T/(2\pi)=10^{-18}\ \mathrm{Hz\cdot K}$, $\rho=21\ \mathrm{g/cm^{3}}$, $g_{0}^{\pm}=0.2$, and $\omega_{\rm fb}^{\pm}=5\Omega_{\pm}$, corresponding to the star symbol in Fig. 2. From these parameters, we obtain $\delta=0.48$ and $\Sigma-\det(\bm{\mathcal{V}})=1.02$. Since the parameters are configured to keep the feedback noise sufficiently low, we can extract the shot noise level, the combined thermal and radiation pressure noise level, and the effective susceptibility by fitting the measurement data to the following expression:
@@
Figure 3: An example of the differential-mode ASD, $S_{q_{-}}^{1/2}$, measured over $10^{7}$ seconds. Blue shows the true values, light blue the estimate, black the measured data. Orange is shot noise, green is radiation pressure noise, magenta is thermal noise, yellow is feedback noise, and red circles show fit results. The inset shows the time-series data $q_{-}$ from 0 to 2000 seconds, with colors matching the main figure. Due to shot noise, the measured signal has an amplitude of about 200, causing the inset’s background to appear completely black.
An example of the differential mode amplitude spectral density (ASD) is shown in Fig. 3. The ASD is obtained using Welch’s method [34], with 50% overlap and a Hanning window. From about 100 simulation runs, we compute the probability density functions of $\Sigma-\det(\bm{\mathcal{V}})$, $\bm{V}_{+}$, and $\bm{V}_{-}$, with four integration times: $10^{5}$ s, $10^{6}$ s, $5\times 10^{6}$ s , and $10^{7}$ s (for details see Appendix D).
@@
where $p(\Sigma-\det(\bm{\mathcal{V}}))$ is the probability density function. For each measurement time, the cumulative probabilities $P_{\text{obs}}^{\text{GIE}}$ are 8.1%, 17.9%, 55.6%, and 72.7%, respectively. As the result demonstrates an improvement proportional to the square root of the measurement time, extrapolating it suggests that the measurement time required to achieve a probability of 99.7% is $1.8\times 10^{7}$ s. Although the measurement time becomes 372 days when the 56% duty cycle is taken into account for 18 mg test masses, it is the same as the one-year operational period of LPF.
@@
For comparison, we calculate that the values of $P_{\text{obs}}^{\text{GIE}}$ are 0%, 3.3%, 15.9%, and 25.9%, respectively, for $\delta=0.28$ (which corresponds to $\Omega_{+}/(2\pi)=0.001$ Hz). In this case, the value of $\Sigma-\det(\bm{\mathcal{V}})$ is 1.007, and compared to the value of 1.02 at $\delta=0.48$, there is a three-fold difference in $\Delta$. Therefore, the integration time required to achieve 99.7% confidence is expected to differ by a factor of 10. The result supports this prediction.
Figure 4: Contour plot of $\log(\Gamma T/(2\pi))$ in the $\log m$-$T_{\rm env}$ plane at $\alpha_{+}=0.081\ \mathrm{Hz^{1/2}}$.
@@
In terms of decoherence, the red region in Fig. 4 shows the region where the total dissipation given by Eqs. (8), (9), and (10) satisfies $\Gamma T/(2\pi)\leq 10^{-18}\ \mathrm{Hz\cdot K}$. Here, we consider a cylindrical mass with 3% absorption at a wavelength of 1064 nm [35, 36], having radius $R=L\sqrt{2/\pi}$ and height $h=L$, which yields $\Lambda=2$ for $m/L^{3}=\rho\Lambda$ [21]. These test masses are surrounded by a shield at temperature $T_{\rm env}$. We assume that the residual gas is an ideal gas consisting of hydrogen at a pressure of $5\times 10^{-16}\ \mathrm{Pa}$ at 4 K. The test masses, with an emissivity of 0.04 [37], are cooled by radiation.
@@
By increasing the mass of the test masses, the gas damping can be mitigated. However, this comes at the cost of increased decoherence because of blackbody radiation emission. This trade-off is important in vacuum, where the levitated mass can only dissipate heat via radiation. As the mass increases, the incident laser power must be increased to maintain the value of $\alpha_{\pm}$, causing radiative cooling to become less effective. Consequently, the equilibrium temperature increases. For large masses, blackbody radiation becomes more influential, whereas for small masses, gas damping becomes more significant. As a result, the optimal mass scale for the GIE is on the order of milligrams. For example, when a 100 mg test mass is placed inside a shield cooled to 2 K and cooled by radiation, it attains an equilibrium temperature of 11 K under an absorbed laser heating power of 0.6 nW. This results in a value of $\Gamma T/(2\pi)=9.3\times 10^{-19}\ \mathrm{Hz\cdot K}$.
In principle, two key conditions must be validated on Earth before realizing the proposed space-based experiment. First, to observe the GIE, the gravitational interaction between the test masses must dominate over all other forces. Since gravitational coupling between milligram-scale objects has already been demonstrated [38], this condition can be satisfied. Second, achieving extremely high vacuum ($<10^{-15}$ Pa) and cryogenic temperatures ($\sim$2 K) in the space environment requires precise engineering design and quantitative validation. Although the required vacuum level is challenging, the ground-based experiment has already achieved $5\times 10^{-15}$ Pa at 4.2 K by cryopomping [39]. For the space mission, a 1 K–class Joule–Thomson cryocooler [40]—with a lifetime exceeding 3 years and a nominal cooling power of 10 mW at 1.7 K—is suitable. A cryocooler capable of cooling down to 0.05 K has also been developed [41]. The minimum value of $\Gamma T/(2\pi)$ is $2.6\times 10^{-19}\ \mathrm{Hz\cdot K}$, achieved at $T_{\mathrm{env}}=0.05\ \mathrm{K}$ and $m=18\ \mathrm{mg}$, as marked by the × in Fig. 4. Since the saturated vapor pressure decreases exponentially as the temperature drops [42, 43], it is preferable to lower the temperature as much as possible. Furthermore, key technical noise sources such as drag-free control, magnetic field effects, and charge-induced noise have already been characterized on LPF [24, 44, 45, 46]. This technological continuity indicates that the challenges faced by the proposed experiment can be reduced by applying systems already demonstrated in space.
Summary— Experimental investigation of the quantum nature of gravity is crucial to advancing modern physics, yet it remains a challenge. In this study, we demonstrate that such an investigation is marginally achievable with current technology. To this end, we simulated quantum measurements of two adjacent mg-scale test masses separated by a mm-scale distance, using a cavity-free interferometer. By improving the vacuum level less than $10^{-15}$ Pa, surrounding the masses with a shield below 2 K, and reducing the resonance frequency $\Omega/(2\pi)$ to $7.7\times 10^{-4}$ Hz, GIE can be demonstrated.
@@
Acknowledgement— We thank Markus Arndt and Ichiro Arakawa for valuable discussions. We thank Koichi Ichimura and Tomoki Fujii for providing the Geant4 source codes. N.M. is supported by JST FORESTO Grant No. JPMJFR202X. K.Y. and A.M. are supported by JSPS KAKENHI Grant No. JP23H01175, and D.M. is supported by JSPS Overseas Research Fellowships.

### v2

# Space-based cm/kg-scale Laser Interferometer for Quantum Gravity

@@ Abstract
The experimental verification of the quantum nature of gravity represents a milestone in quantum gravity research. Recently, interest has grown for testing it via gravitationally induced entanglement (GIE). Here, we propose a space-based interferometer inspired by the LISA Pathfinder (LPF). Our design employs two kg-scale gold-platinum test masses which, unlike in the LPF, are surrounded by a shield below 1 K and positioned side-by-side with a centimeter-scale separation. This configuration enables the detection of GIE through simultaneous measurements of differential and common-mode motions. To estimate the integration time required for GIE detection, we simulate quantum measurements of these modes, considering noise sources such as gas damping, black-body radiation, and cosmic-ray collisions. Our results show that GIE can be demonstrated with a few modifications to the LPF setup.
@@
Here, we propose a cavity-free laser interferometer that optimizes both signal amplification and direct noise reduction in optomechanical systems, in contrast to previous cavity-based studies [19, 21, 14, 17, 23]. Our proposed experiment closely resembles the space-based interferometer LISA Pathfinder (LPF) [24], which marked the beginning of a new era in interferometric experiments; however, it differs in the following three key aspects: (1) The test masses are placed close to enable the generation of gravitationally induced entanglement (GIE); (2) Both differential and common-mode motions are measured to enable the detection of GIE via Kalman filtering; and (3) The masses are surrounded by a shield below 1 K, and both modes of motion are further cooled via feedback using a high-pass filter. Using the space environment, the resonance frequency of the test mass can be reduced to 0.25 mHz, effectively amplifying the gravitational coupling signal while remaining free from ground vibrations, thereby making it possible to observe the GIE within a total integration time of about 40 days.
Figure 1: Experimental setup. The differential and common-mode of test mass 1 (TM1) and test mass 2 (TM2) are measured by laser interferometry with a wavelength of 1064nm1064\ $\mathrm${nm}.
@@
where $Y_{\pm}$ represent the phase quadratures, $\bm{C}_{\pm}=(-\alpha_{\pm},0,0)$ denote the optomechanical coupling constants, and $\bm{x}_{\pm}=(q_{\pm},p_{\pm},r_{\pm})^{T}$ denote the position, momentum, and auxiliary variables used to describe feedback cooling (for details see Appendix A). The terms $y^{\rm in}_{\pm}$ represent vacuum (white) noise, with variances given by $\langle(y_{\pm}^{\rm in})^{2}\rangle=1$. In addition, we introduce vacuum noise $x^{\rm in}_{\pm}$ for the amplitude quadratures. Throughout this work, the subscripts ++ and - refer to the common mode and the differential mode, respectively. The optomechanical coupling constants $\alpha_{\pm}$ are given by $\sqrt{16\omega_{c}P_{\rm in}/(m\Omega_{\pm}c^{2}})$, where $\omega_{c}$ is the laser frequency and $P_{\rm in}$ the incident laser power, $m$ the mass of each mirror, $\Omega_{+}=\Omega$ and $\Omega_{-}=\Omega\sqrt{1-\delta}$ the resonance frequency of the common and differential modes. The term $\delta=\Lambda\times 4Gm/(\Omega^{2}L^{3})$ represents the gravitational coupling between the two mirrors, where $\Omega$ is the resonance frequency of each mirror, $L$ is the average distance between their centers of mass, and $\Lambda$ is a form factor that accounts for the deviation from a point-mass approximation (see Appendix B).
@@
Here, $\bm{\mathcal{V}}$ is the conditional covariance matrix of the individual test masses (see Appendix C) and $\Sigma=\det\bm{\mathcal{V}}_{1}+\det\bm{\mathcal{V}}_{2}-2\det\bm{\mathcal{V}}_{12}$, where $\bm{\mathcal{V}}_{1}$, $\bm{\mathcal{V}}_{2}$ are the covariance matrices of TM1 and TM2 normalized by the frequency $\Omega$, respectively. $\bm{\mathcal{V}}_{12}$ represents the gravity-induced correlation matrix between individual masses. According to the separability condition for two-mode Gaussian states, the systems are entangled if and only if $E_{N}>0$ [27]. To verify whether quantum entanglement is present from Eq. (5), it is sufficient to confirm $\Sigma-\text{det}\bm{\mathcal{V}}>1$ for $\Sigma>2$.
The colored region in Fig. 2 shows the region where GIE is generated. The generation rate derived in Ref. [18] must be faster than the thermal decoherence rate, as shown below:
@@
From the above inequality, the condition for the GIE is bounded by
${\left({\Gamma T/2\pi\over 6.5\times 10^{-19}{\ \rm Hz\cdot K}}\right)}{\left({20000{\ \rm kg/m^{3}}\over\rho}\right)}\left({1\over\Lambda}\right)<1,$ (7)
where $T$ represents the temperature of the test masses and $\rho$ is the mass density of the mirror.
Figure 2: Contour plot of $E_{N}$ in the $\alpha_{+}$-$\Gamma T/(2\pi)$ plane. The green dashed line indicates the threshold for generating the GIE in Eq. (6), the red curve shows the boundary where the relaxation time due to the Kalman filter is equal to half the thermal decoherence time (see Appendix A). The star symbol corresponds to the parameters used in the simulation. The GIE can be generated when $\Gamma T/(2\pi)<4.5\times 10^{-19}$ Hz K.
@@
Finally, we simulate the acceleration noise from cosmic-ray collisions using Geant4 [30] and a flux model [31][32]. After removing large events through data processing, the requirement is met with a duty cycle—the fraction of usable data time—of 12% (see Appendix D).
@@
In the simulation, we set $\Omega_{+}/(2\pi)=2.5\times 10^{-4}\ \mathrm{Hz}$, $\alpha_{+}=0.031\ \mathrm{Hz^{1/2}}$, $\Gamma T/(2\pi)=2\times 10^{-19}\ \mathrm{Hz\cdot K}$, $\rho=20\ \mathrm{g/cm^{3}}$, $g_{0}^{+}=0.2$, $\omega_{\rm fb}^{+}=5\Omega_{+}$, $g_{0}^{-}=1$, and $\omega_{\rm fb}^{-}=2\Omega_{-}$. For a cylindrical test mass with thickness $h$, radius $R=0.5h$, and center-of-mass separation $L=1.25h$, we find that $\delta=0.91$ and $\Sigma-\det(\bm{\mathcal{V}})=1.27$. For a 1 kg test mass, the zero-point amplitude is $q^{+}_{\rm zpf}=1.8\times 10^{-16}$ m and the surface-to-surface gap is $h/4\simeq 1$ cm. The required resonant frequency of $\Omega_{+}/(2\pi)=2.5\times 10^{-4}\ \mathrm{Hz}$ can be achieved either passively, through the gravitational field of ancillary tuning masses, or actively, using an electrostatic feedback system.
Since the above parameters are configured to keep the feedback noise sufficiently low, we can extract the shot noise level, the combined thermal and radiation pressure noise level, and the effective susceptibility by fitting the measurement data to the following expression:
@@
Figure 3: An example of the common-mode ASD $S_{q_{+}}^{1/2}$, normalized by the zero-point amplitude $q^{+}_{\rm zpf}$, measured over $3\times 10^{6}$ seconds. For a 1 kg test mass, the value 1 on the vertical axis corresponds to $1.8\times 10^{-16}\ {\rm m/\sqrt{Hz}}$. Blue shows the true values, light blue the estimate, black the measured data. Yellow is shot noise, green is radiation pressure noise, magenta is thermal noise, orange is feedback noise, and red circles show fit results. The inset shows the time-series data $q_{+}$ from 0 to $2\times 10^{4}$ seconds, with colors matching the main figure. Due to shot noise, the measured signal has an amplitude of about 20, causing the inset’s background to appear black.
An example of the common mode amplitude spectral density (ASD) is shown in Fig. 3. The ASD is obtained using Welch’s method [35], with 50% overlap and a Hanning window. From about 100 simulation runs, we compute the probability density functions of $\Sigma-\det(\bm{\mathcal{V}})$, $\bm{V}_{+}$, and $\bm{V}_{-}$, with four integration times: $10^{5}$ s, $5\times 10^{5}$ s, $2\times 10^{6}$ s , and $3\times 10^{6}$ s (for details see Appendix E).
@@
where $p(\Sigma-\det(\bm{\mathcal{V}}))$ is the probability density function. For each measurement time, the cumulative probabilities $P_{\text{obs}}^{\text{GIE}}$ are 13%, 34%, 79%, and 88%, respectively. As the result demonstrates an improvement proportional to the square root of the measurement time, extrapolating it suggests that the measurement time required to achieve a probability of 99.7% is $3.5\times 10^{6}$ s. Although the measurement time becomes one year when the duty cycle of 10% is taken into account for 1 kg test masses, it is the same as the operational period of LPF.
@@
For comparison, we calculate the values of $P_{\text{obs}}^{\text{GIE}}$, for $\delta=0.75$, which corresponds to the distance of $L=4h/3$. In this case, the value of $\Sigma-\det(\bm{\mathcal{V}})$ is 1.09, and compared to the value of 1.27 at $\delta=0.91$, there is a 3-fold difference in $\Delta$. Therefore, the integration time required to achieve 99.7% confidence is expected to differ by a factor of $\sim 10$ (see Appendix E). The result supports this prediction.
Figure 4: Contour plot of $\log(\Gamma T/(2\pi))$ in the $\log m$-$T_{\rm env}$ plane at $\alpha_{+}=0.031\ \mathrm{Hz^{1/2}}$.
@@
In terms of decoherence, the red region in Fig. 4 shows the region where the total dissipation given by Eqs. (8), (9), and (10) satisfies $\Gamma T/(2\pi)\leq 4.5\times 10^{-19}\ \mathrm{Hz\cdot K}$. Here, we consider a cylindrical mass with 3% absorption at a wavelength of 1064 nm [36, 37], with radius $R=0.5h$ and center-of-mass separation $L=1.25h$. These test masses are surrounded by a shield at temperature $T_{\rm env}$. We assume that the residual gas is an ideal gas consisting of hydrogen at a pressure of $5\times 10^{-15}\ \mathrm{Pa}$ at 4 K. The test masses, with an emissivity of 0.04 [38], are cooled by radiation.
@@
By increasing the mass of the test masses, the gas damping can be mitigated. However, this comes at the cost of increased decoherence because of blackbody radiation emission. This trade-off is important in vacuum, where the levitated mass can only dissipate heat via radiation. As the mass increases, the incident laser power must be increased to maintain the value of $\alpha_{\pm}$, causing radiative cooling to become less effective. Consequently, the equilibrium temperature increases. For large masses, blackbody radiation becomes more influential, whereas for small masses, gas damping becomes more significant. As a result, the optimal mass scale for the GIE is on the order of grams to kilograms. For example, when a 1 kg test mass is placed inside a shield cooled to $T_{\rm env}=0.1$ K and cooled by radiation, it attains an equilibrium temperature of $T=11.7$ K under an absorbed laser heating power of $2\times 0.14$ uW. This results in a value of $\Gamma T/(2\pi)=2.0\times 10^{-19}\ \mathrm{Hz\cdot K}$.
In principle, three key conditions must be validated before realizing the proposed space‐based experiment. First, achieving extremely high vacuum ($5\times 10^{-15}$ Pa) and cryogenic temperatures ($\sim$0.1 K) in the space environment requires precise engineering design and quantitative validation. Although the required vacuum level is challenging, the ground-based experiment has already achieved $5\times 10^{-15}$ Pa at 4.2 K by cryopumping [39]. For space missions, a cryocooler capable of cooling down to 0.05 K has been developed [40]. This Adiabatic Demagnetization Refrigerator (ADR) provides a nominal cooling power of 0.4 µW at 0.05 K, which is greater than the heat input from the laser. A spherical cryogenic shield (e.g.,  20 cm inner diameter,  50 kg) would by itself create a resonant frequency of 0.29 mHz. To counteract this and passively set the frequency to within a few percent of the target, two external tuning masses (e.g., 88 kg each at 30 cm) provide a counteracting gravitational gradient. Any residual offsets can then be precisely corrected using electrostatic feedback. Importantly, a custom launch-lock mechanism is required to operate in the cryogenic, ultra-high vacuum environment.
Second, unlike LPF, this study must precisely measure not only the differential motion of the test masses but also their common-mode motion. In LPF, the spacecraft is controlled with respect to one test mass (drag-free control) to mitigate disturbances such as the solar wind; however, the achievable stability is limited by noise from the thrusters used for control, which leads to elevated noise levels in the common-mode motion. To avoid this, noise reduction via passive shielding is effective (for details see Appendix F).
Finally, acceleration noise from patch effects remains a significant challenge that requires further investigation, particularly at cryogenic temperatures and low frequencies. This noise arises from electrostatic forces between adjacent metallic surfaces—such as those between test masses and actuators—and is driven by fluctuations in both charge and surface patch potentials. Its behavior is frequency-dependent: high-frequency components have been observed to decrease significantly under cryogenic conditions [41], whereas its low-frequency characteristics are not yet well characterized.
Effective mitigation strategies include general approaches such as shielding and charge neutralization. Moreover, the noise level can be further reduced through careful design, as the patch force itself decreases with greater surface separation, while the resulting acceleration is inversely proportional to the test mass. Our design takes advantage of both principles by incorporating a large (1 kg) test mass and a wide (1 cm) separation between surfaces, and is therefore expected to be highly effective in suppressing this noise source.
Summary— Experimental investigation of the quantum nature of gravity is crucial for advancing modern physics, yet it remains a significant challenge. In this study, we show that such an investigation is marginally achievable with current technology. To this end, we simulated quantum measurements of two adjacent kg-scale test masses separated by a cm-scale distance. Our simulation results indicate that a detectable signature—Gravitationally Induced Entanglement (GIE)—can be generated between the masses. Achieving an experimental demonstration of this GIE, however, requires stringent conditions. These include attaining the highest vacuum levels possible in terrestrial experiments, cooling the test masses inside state-of-the-art refrigerator. Although a deeper understanding of low-temperature, low-frequency patch potentials is necessary to demonstrate the GIE, mitigating this noise source appears feasible, suggesting that experimental observation of GIE is within reach.
@@
Acknowledgement— We thank Markus Arndt and Ichiro Arakawa for valuable discussions, and Koichi Ichimura and Tomoki Fujii for providing the Geant4 source codes. N.M. conceived and designed the research, performed simulations with K.S., and wrote the manuscript. K.S. developed the state-space model, and K.H. carried out the GEANT4 simulations. All authors discussed the results. N.M. is supported by JST FOREST Grant No. JPMJFR202X. K.S. is supported by RIKEN TRIP initiative (RIKEN Quantum). K.Y. and A.M. are supported by JSPS KAKENHI Grant No. JP23H01175, and D.M. is supported by JSPS Overseas Research Fellowships.